OpenStudy (anonymous):

Robyn rolled a number cube 60 times in an experiment. Her results are shown in this table. Number Rolled= 1 2 3 4 5 6 Number of Times= 8 5 9 12 11 15 What is the experimental probability of rolling a 6? A 25% B 40% C 60% D 10% Thanks for anyone who can assist me with this problem. =)

OpenStudy (lucaz):

the cube was rolled 60 times and he got 6 15 times P(6) = # times / # total = 15/60=1/4 or (1/4)*100=25%

OpenStudy (anonymous):

so would the correct answer be a or c? im stuck on those...thnx

OpenStudy (lucaz):

A, its 25%

OpenStudy (anonymous):

ty very much